Find the LCM and HCF of the following pairs of integers and verify that LCM × HCF = product of the two numbers.
336 and 54

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Expressing 336 and 54 as product of its prime factors, we get,
336 = 2 × 2 × 2 × 2 × 7 × 3 × 1
54 = 2×3×3×3×1
Therefore, LCM(336, 54) =2^4× 3^3× 7= 3024
And HCF(336, 54) = 2×3 = 6
Verification
Now, product of 336 and 54 = 336 x 54 = 18144
And Product of LCM and HCF =3024 ×4 = 18144
Hence,LCM × HCF = product of the 336 and 54.

​Thanks